An alert physics student stands beside the tracks as a train rolls slowly past. He notes that the frequency of the train whistle is 485 Hz when the train is approaching him and 458 Hz when the train is receding from him. Using these frequencies, he calculates the speed of the train. What value does he find? (Assume the speed of sound in air is 343 m/s.) m/s
Answer #1

The apparent frequency f' = f *c/(c -v) --> 485 = f*343/(343-v).........................(1)
f" = f*c/(c+v) ---> 458 = f *343/(343+v) ..................................(2)

-Divide equation (1) and (2)

485/458 = (f*343 *(343+v))/ (f *343*(343-v))

v = 9261/943 = 9.82 m/s